OpenStudy (anonymous):

How do I find the inverse function for this doubling function: y=100(2)^(x/0.32)? Can someone please show me? :'(

OpenStudy (anonymous):

This is what I have so far:

OpenStudy (anonymous):

y= 100(2)^(x/0.32) x = 100(2)^(y/0.32) x/100 = 2^(y/0.32) logx/100 = (y/0.32)log2 0.32logx/100 = ylog2 (0.32logx/100)/2 = y HELP! D; The answer is supposed to be y= 0.32log of base 2(x/100) I don't know where I'm going wrong ;'( @Hero and @TuringTest do you guys know how to do this?

hero (hero):

It's not difficult

OpenStudy (anonymous):

really? ahh, can you please show me how?

hero (hero):

Yeah you made a mistake

OpenStudy (anonymous):

Your last step was wrong. I used natural logs rather than log10, but when you get to 0.32[ln(x/100)/ln2] it becomes what you were supposed to get. You seemed to have just removed the log from the 2.

hero (hero):

y= 100(2)^(x/0.32) x = 100(2)^(y/0.32) x/100 = 2^(y/0.32) log(x/100) = (y/0.32)log(2) log(x/100)/log(2) = y/0.32 (0.32(log(x/100)/log(2))) = y

hero (hero):

Parentheses are a little off, but yeah, that's it

OpenStudy (anonymous):

wait oops I should have put the log2, anyway the answer is:

OpenStudy (anonymous):

so wait why does it say log of base 2... not log of base 10

hero (hero):

Should be log[10](2)

hero (hero):

I think the answer is equivalent anyways

OpenStudy (anonymous):

really? ...? o.O not: \[y=0.32log _{2}(x/100)\]

hero (hero):

I'm pretty sure it is equivalent

OpenStudy (anonymous):

ohh.. okay thanks!

hero (hero):

They must have used some change of base formula to simplify

OpenStudy (anonymous):

or could it be because: log[a](b) = log b/log a and so for this situation, [0.32log(x/100)]/log2 --> 0.32log[2](x/100)? since log2 could be the log a and log(x/100) could be log b?

hero (hero):

they used the change of base formula bro